# Shear flow of frictional spheroids: lentil-like grains slide more than rice-like ones

### v1

@@ III.3 Rheological properties
Figure 7: Stress anisotropy (top) and packing fraction (middle), and average number of contacts per particle (bottom) for different values of microscopic friction. The dashed, dotted-dashed and dotted lines indicate $Z=3,6,10$ respectively. Left $\mu_{p}=0$, middle $\mu_{p}=0.4$, right $\mu_{p}=10.0$. Same color legend as in Figure 6.
Having characterized the fluctuations of the grains, we move on to their rheology. In this section we present the best fit of effective friction and packing fraction with respect to the standard $\mu(I)$ rheology [55], $\mu(I)=\sigma_{xy}/P$, for different miscroscopic friction, see Figure 7a and b. The well-known formulas are:
@@
where $\mu_{c},\,\phi_{c}$ are the critical effective friction and packing fraction respectively, while $I_{0},c_{\mu},c_{\phi}$ are constants and $\beta_{\phi}$ is an exponent. Similarly to [21, 22] the effective friction generally increases by increasing particle elongation (or flatness) and microscopic friction. The actual values for $\alpha=2$ are almost identical to those measured by [22] for spherocylinders with the same aspect ratio, suggesting that $\alpha$ is a good parameter to characterize the shape in granular rheology, beyond the characteristic size of the grain. See Supplemetary Figure 1 for a direct comparison between spheroids and data for spheroids taken from [22].
We observe that it is slightly easier to shear flat particles for the frictionless case with respect to elongated ones with the same value of $|r_{g}|$, which is also reflected by the fact that oblate spheroids attain a higher alignment, which minimizes $A_{yz}$. For $\mu_{p}=0.4$ slightly oblate particles, $0.66\leq\alpha\leq 1.0$, are harder to shear, but the opposite is true for more elongated ones. As $\mu_{p}$ increases, deviations from the spherical shape make shearing harder, since the alignment that made sliding efficient without friction, now comes at a higher cost. Nevertheless as the mechanism of dissipation reverts back to collisional at very high microscopic friction, the effective friction of the granular assembly keeps a finite values, even in the limit $\mu_{p}\to\infty$.
@@
Similarly to the shear strength, the effect of friction on the packing fraction is highly nonlinear, see Figures 7b and 8b. Elongated particles reach higher densities in the low friction limit, whereas the opposite is true as friction comes into play. This behaviour is similar to what is observed for random packings of non-spherical particles [56, 57], with $\mu_{p}$ providing the randomizing effect. Interestingly, the values of $\phi_{c}$ for low friction and small eccentricity lie close to those measured by [56] and defined as random close packings. However the effect of shear becomes prominent at high degree of eccentricity where grains can achieve significantly higher density than if randomly packed.
Figure 8: Critical shear strength (top) and packing fraction (bottom) as a function of aspect ratio. For $\phi_{c}$ values of random close packing from [56] are also reported with a black line and open circles, while values for $\mu_{p}=10.0$ are not reported since simulations at the lower values of $I$ are needed to get a meaningful estimate. Same color legend as in Figure 2.
@@
For high friction, aspherical particles achieve a lower volumetric density and their coordination becomes similar to that of spheres. Additional results on normal stress differences can be found in Appendix D.
@@ III.4 Correlation of velocity fluctuations
As stated earlier, the higher correlation of motion for flattened particles, is another factor leading to the extended sliding regime they display compared to elongated ones, since to break the correlation higher values of $\mu_{p}$ or $I$ are needed. To better characterize the correlation length, especially in the low-inertial-number regime where it tends to diverge, additional simulations with larger system sizes are required to exclude finite size effects. Nevertheless, the values reported here are reasonably smaller than the system size, with the maximum ratio of $\ell/L_{y}=0.16$ measured at $I=0.1$, and the trend shown is consistent along all inertial numbers object of this study: although not quantitatively rigorous, the qualitative trends observed are deemed trustworthy.
@@ III.5 Contact fabric in the sliding regime
Figure 10: Fabric eigenvectors scaled by their corresponding eigenvalue at $I\approx 0.01$. In the inset: the angle $\eta$ (in blue) between the largest eigenvectors of the reciprocal shapes, e.g. $\alpha_{1}=3,\,\alpha_{2}=1/\alpha_{1}=1/3$, the ratio between the biggest eigenvalue of the prolate spheroid with respect to the reciprocal oblate one, $\lambda_{1}^{p}/\lambda_{1}^{o}$, in red. Same color legend as in Figure 6.
In Figure 10 we report the two eigenvectors in the $xy$ plane scaled by their respective eigenvalue. For spheres the eigenvectors are roughly oriented at ±$\pm\pi/4$, in the extension and compression direction of the flow respectively. As $|r_{g}|$ increases, the eigenvectors rotate clockwise similarly to the force distribution, see Section III.5. The eigenvalue in the compression direction is greater than in the extension one in the spherical case. The anisotropy measure as the ratio of the two biggest eigenvalues increases with eccentricity. Additionally, one notices that the eigenvalues in both the compression and extension direction of oblate particles are greater than the one for prolate ones, suggesting that fewer contacts take place in the vorticity direction for lentil-like particles. The full planar distribution of contacts can be found in Appendix F. As shown in the inset of Figure 10, the anisotropy of oblate particle is higher than the corresponding reciprocal prolate ones. Despite alignment of reciprocal shapes being similar, recall Figure 2a, the much more anisotropic fabric makes this alignment harder to disrupt, hence the transition to the rolling regime happens for higher values of $\mu_{p}$. The angle of the principal eigenvectors shows only a modest discrepancy.
@@ III.6 Particle contact data in the sliding regime
[…]
Figure 11: Polar histograms for contact data in the ellipsoid reference frame. The angle is measured from the ellipsoid symmetry axis. (a) Surface probability distribution of contacts, $\rho_{s}$. (b, c) Average normal and tangential force conditional to a contact being present normalized by average pressure times the surface area of the particle, $\overline{F}_{n}/PA_{p}$, and $\overline{F}_{t}/PA_{p}$ respectively. It appears evident that the contact density and the force intensity are anticorrelated. Same color legend as in Figure 6.
@@
Next, we look at the local friction mobilization defined as the ratio $\mathcal{M}=F_{t}/\mu_{p}F_{n}$. A value of $\mathcal{M}=0$ means that all tangential springs are inactive, whereas a value of $\mathcal{M}=1$ indicates that all contacts have reached the sliding limit. In Figure 12a we report data for $\mu_{p}=0.4$ and different degrees of elongation. As found in previous studies [59, 23], elongated particles mobilize more friction compared to spherical ones. The higher mobilization is closely related to the extended frictional sliding regime shown in Figure 3. The cause is that spherical particles can more easily rotate when a tangential force is applied, contrary to non-spherical ones (see Section III.3). Interestingly, for low degrees of eccentricity oblate particles mobilize more friction than prolate ones, while for higher eccentricity the converse is true. This nonlinearity reflects the same nonlinearity observed for the effective friction $\mu(I)$, due to the ordering.
Defining $\mathcal{P}_{j}$ as the average power dissipated on the surface area element $A_{j}$ conditional to a contact being present, and $\mathcal{P}_{p}$ as the total average power dissipated on the particle surface, we compute the spatial inhomogeneity of power dissipation density as $\mathcal{I}=\mathcal{P}_{j}\rho_{s}A_{p}^{2}/(\mathcal{P}_{tot}A_{j})$. A value of $\mathcal{I}=1$ indicated a homogeneous dissipation, while values $>1$ and $<1$ indicate hot spots and cold spots, respectively. It can be observed, see Figure 12b, that the local power dissipation on the surface of the particles clusters in the same locations, where slip events are more common. For high degrees of eccentricity, more than ten times dissipation occurs on the major axis compared to the minor one. Accordingly, we propose the following microscopic mechanism: non-spherical particles establish a small number of strong contacts along their minor axis, where sliding is rare. To accommodate the linear velocity gradient, the particles exhibit a wobbling motion, resulting in the majority of power dissipation occurring at weaker contacts along the major axis. A more detailed microscopic analysis of the contribution of frictional forces to shear strength can be found in Appendix G.
@@ IV Summary and outlook
Our regime analysis revealed that flat, lentil-like particles exhibit an extended sliding regime compared to elongated, rice-like particles, even when the shape ratio is kept constant. This behavior was further clarified by examining the fluctuating energy components: oblate and prolate particles differ in how this energy is distributed and expressed in their dynamics. Moreover, slightly aspherical particles at low friction exhibit non-monotonic behavior, which we linked to enhanced rotational kinetic energy fluctuations.
@@
This work can be extended in several directions. First, the relationship between fabric anisotropy and stress transmission in non-spherical particle assemblies could be further examined within the framework proposed by Srivastava et al. [60]. Additionally, incorporating frictional effects into the anisotropic rheological model proposed by Nadler et al. [25] could enhance predictive capabilities for both natural and industrial applications.
Another key avenue for investigation is the role of vortices and clusters in shear-driven particulate flows, as discussed in [54, 37, 61]. The formation and dynamics of vortical structures could significantly impact shear banding and localized dissipation, necessitating a deeper physical understanding. Moreover, extending our analysis to dense suspensions [29] will help bridge the gap between granular and suspension rheology, elucidating the transition between friction-dominated and hydrodynamic-dominated flow regimes. Lubrication forces may become dominant at lower viscous numbers, especially for flat particles which enhance lubrication effects. Finally, a systematic study should be conducted on non-convex, potentially interlocking particles [62]. A deeper understanding of these interactions could inform the design of polycatenated materials [63] with novel particle architectures, ultimately improving control over their flow properties.
@@ Acknowledgements
We gratefully thank Professor Ali Ozel for sharing the preliminary input scripts on simple shear under quasi-static conditions, which contributed significantly to the work presented in [64]. We also acknowledge Lucas Fourel for expert assistance with visualization software and insightful advice on data presentation. This project was supported by the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 945363. We further appreciate the financial support from Bühler AG and the valuable discussions with Clément Zemerli and Richard Vonlanthen.
@@
Data to support these finds is available on Zenodo [65] and will be made public upon publication. Codes to run the simulations and the analysis will be made available on GitHub.
@@ Appendix A Contact interaction
$T_{H}=$ $2.94\left(5\sqrt{2}\pi\rho_{p}\over 4E^{*}\right)^{2/5}{a_{min}\over\dot{\gamma}d_{eq}}\,$ (23)
$T_{R}=$ ${\pi a_{min}\over 0.8766+0.163\nu}\sqrt{\rho_{p}\over G}\,$ (24)
where $G$ is the shear modulus and we assumed the typical impact velocity to be of the order of the one imposed by the shear flow ($\dot{\gamma}d_{eq}$).
@@
subject to $\text{subject to }G_{1}(\mathbf{x})=G_{2}(\mathbf{x})\,.$ (26)
@@ Appendix B Dissipation computation
$\mathcal{P}_{n}$ $=\sum_{N_{c}}\left[\frac{\mathbf{F}^{\gamma}_{n,i}}{2}\cdot\left(\mathbf{v}_{n,i}+\frac{1}{4m*}\mathbf{F}_{n,i}\right)+\right.$
$\quad\left.\frac{\mathbf{F}^{\gamma}_{n,i+1}}{2}\cdot\left(\mathbf{v}_{n,i+1/2}+\frac{1}{4m*}\mathbf{F}_{n,i+1}\right)\right]$
$\approx\sum_{N_{c}}\mathbf{F}_{t,i}\cdot\mathbf{v}_{t,i}\,,$ (27)
@@
$\mathcal{P}_{n}=\sum_{N_{s}}\mathbf{F}_{t,i}\cdot\left(\frac{\bm{\delta}_{t,i}+\mathbf{F}_{t,i}/k_{t}}{\Delta t}\right)\approx\sum_{N_{s}}\mathbf{F}_{t,i}\cdot\mathbf{v}_{t,i}$ (28)
where the second equality is satisfied for small enough $\Delta t$ so that slip only occurs at the beginning of the time step and the tangential velocity can be considered constant during the interval. We verified both Equations 27 and 28 this by halving the time step size. In particular Equation 27 is tested for those simulations with the highest fluctuations $\mu_{p}=10$, $I=0.1$, corresponding to dominant collisional dissipation, while Equation 28 is tested for $\mu_{p}=0.4$, $I=0.01$, where frictional dissipation is dominant. Figure 13 shows results within statistical error, implying that the chosen time step is sufficiently small.

### v2

@@ III.3 Rheological properties
Figure 7: Stress anisotropy (top) and packing fraction (middle), and average number of contacts per particle (bottom) for different values of microscopic friction. The dashed, dotted-dashed and dotted lines indicate $Z=3,6,10$ respectively.
Having characterized the fluctuations of the grains, we move on to their rheology. In this section we present the best fit of effective friction and packing fraction with respect to the standard $\mu(I)$ rheology [63], $\mu(I)=\sigma_{xy}/P$, for different miscroscopic friction, see Figure 7a and b. While more fundamental kinetic theories exist, this phenomenological model is chosen for its proven effectiveness in capturing the essential flow behavior and its role as a standard benchmark, allowing for direct comparison with other studies in the field. The well-known formulas are:
@@
where $\mu_{c},\,\phi_{c}$ are the critical effective friction and packing fraction respectively, while $I_{0},c_{\mu},c_{\phi}$ are constants and $\beta_{\phi}$ is an exponent. Similarly to [21, 22] the effective friction generally increases by increasing particle elongation (or flatness) and microscopic friction. The actual values for $\alpha=2$ are almost identical to those measured by [22] for spherocylinders with the same aspect ratio, suggesting that $\alpha$ is a good parameter to characterize the shape in granular rheology, beyond the characteristic size of the grain. See supplemetary Figure [64] [64] for a direct comparison between spheroids and data for sphero-cylinders taken from [22].
We observe that it is slightly easier to shear flat particles for the frictionless case with respect to elongated ones with the same value of $|r_{g}|$, which is also reflected by the fact that oblate spheroids attain a higher alignment, which minimizes $A_{yz}$. For $\mu_{p}=0.4$ slightly oblate particles, $0.66\leq\alpha\leq 1.0$, are harder to shear, but the opposite is true for more elongated ones. As $\mu_{p}$ increases, deviations from the spherical shape make shearing harder, since the alignment that made sliding efficient without friction, now comes at a higher cost. Nevertheless as the mechanism of dissipation reverts back to normal dissipation at very high microscopic friction, the effective friction of the granular assembly keeps a finite values, even in the limit $\mu_{p}\to\infty$.
@@
Similarly to the shear strength, the effect of friction on the packing fraction is highly nonlinear, see Figures 7b and 8b. In the low friction limit, more elongated or flat particles reach higher densities, whereas these trends become non-monotonic as friction comes into play. This behaviour is similar to what is observed for random packings of non-spherical particles [65, 66], with $\mu_{p}$ providing the randomizing effect. Interestingly, the values of $\phi_{c}$ for low friction and small eccentricity lie close to those measured by [65] and defined as random close packings. However the effect of shear becomes prominent at high degree of eccentricity where grains can achieve significantly higher density than if randomly packed.
Figure 8: Critical shear strength (top) and packing fraction (bottom) as a function of aspect ratio. For $\phi_{c}$ values of random close packing from [65] are also reported with a black line and open circles, while values for $\mu_{p}=10.0$ are not reported since simulations at the lower values of II are needed to get a meaningful estimate.
@@
For high friction, aspherical particles achieve a lower volumetric density and their coordination becomes similar to that of spheres.
The values measured in Figure 7b and c are in most cases slightly higher than the critical values for the appearance of elastic stresses, proportional to $E$, as measured under constant volume conditions in [29], provided a direct comparison between cylinders and spheroids of the same aspect ratio is appropriate. Additional evidence is found on the extended plateaus of $\phi$, $Z$ with respect to inertial number in Figure 7b.2 and 7c.2, which signal a quasi-static (elastic) response [31]. Additional results on normal stress differences can be found in Appendix D.
@@ III.4 Correlation of velocity fluctuations
As stated earlier, the higher correlation of motion for flattened particles, is another factor leading to the extended sliding regime they display compared to elongated ones, since to break the correlation higher values of $\mu_{p}$ or II are needed. To better characterize the correlation length, additional simulations with larger system sizes are required to exclude finite size effects. Notwithstanding, when particles are allowed to overlap in 3D the correlation length is expected to increase up to the volume fraction where shear jamming is reached [68, 69]. As previously mentioned, in several of the simulations, especially those in the sliding regime at low $I$, we are close to this limit. The values reported here are reasonably smaller than the system size, with the maximum ratio of $\ell/L_{y}=0.16$ measured at $I=0.1$, and the trend shown is consistent along all inertial numbers object of this study: although not quantitatively rigorous, the qualitative trends observed are deemed trustworthy.
@@ III.5 Contact fabric in the sliding regime
Figure 10: Fabric eigenvectors scaled by their corresponding eigenvalue at $I\approx 0.01$. In the inset: the angle $\eta$ (in blue) between the largest eigenvectors of the reciprocal shapes, e.g. $\alpha_{1}=3,\,\alpha_{2}=1/\alpha_{1}=1/3$, the ratio between the biggest eigenvalue of the prolate spheroid with respect to the reciprocal oblate one, $\lambda_{1}^{p}/\lambda_{1}^{o}$, in red.
In Figure 10 we report the two eigenvectors in the $xy$ plane scaled by their respective eigenvalue. For spheres the eigenvectors are roughly oriented at $\pm\pi/4$, in the extension and compression direction of the flow respectively. As $|r_{g}|$ increases, the eigenvectors rotate clockwise similarly to the force distribution, see Section F. The eigenvalue in the compression direction is greater than in the extension one in the spherical case. The anisotropy, measured as the ratio of the two biggest eigenvalues, increases with eccentricity. Additionally, one notices that the eigenvalues in both the compression and extension direction of oblate particles are greater than the one for prolate ones, suggesting that fewer contacts take place in the vorticity direction for lentil-like particles. The full planar distribution of contacts can be found in Appendix F. As shown in the inset of Figure 10, the anisotropy of oblate particle is higher than the corresponding reciprocal prolate ones. Despite alignment of reciprocal shapes being similar, recall Figure 2a, the much more anisotropic fabric makes this alignment harder to disrupt, hence the transition to the rolling regime happens for higher values of $\mu_{p}$. The angle of the principal eigenvectors shows only a modest discrepancy.
@@ III.6 Particle contact data in the sliding regime
$\int_{A}\rho_{s}dA=1\,.$ (10)
Figure 11: Polar histograms for contact data in the ellipsoid reference frame. The angle is measured from the ellipsoid symmetry axis. (a) Surface probability distribution of contacts, $\rho_{s}$. (b, c) Average normal and tangential force conditional to a contact being present normalized by average pressure times the surface area of the particle, $\overline{F}_{n}/PA_{p}$, and $\overline{F}_{t}/PA_{p}$ respectively. It appears evident that the contact density and the force intensity are anticorrelated.
@@
Next, we look at the local friction mobilization defined as the ratio $\mathcal{M}=F_{t}/\mu_{p}F_{n}$. A value of $\mathcal{M}=0$ means that all tangential springs are inactive, whereas a value of $\mathcal{M}=1$ indicates that all contacts have reached the sliding limit. In Figure 12a we report data for $\mu_{p}=0.4$ and different degrees of elongation. As found in previous studies [70, 23], elongated particles mobilize more friction compared to spherical ones. The higher mobilization is closely related to the extended frictional sliding regime shown in Figure 3. The cause is that spherical particles can more easily rotate when a tangential force is applied, contrary to non-spherical ones (see Section III.1. Interestingly, for low degrees of eccentricity oblate particles mobilize more friction than prolate ones, while for higher eccentricity the converse is true. This nonlinearity reflects the same nonlinearity observed for the effective friction $\mu(I)$, due to the ordering.
Defining $\mathcal{P}_{j}$ as the average power dissipated on the surface area element $A_{j}$ conditional to a contact being present, and $\mathcal{P}_{p}$ as the total average power dissipated on the particle surface, we compute the spatial inhomogeneity of power dissipation density as $\mathcal{I}=\mathcal{P}_{j}\rho_{s}A_{p}^{2}/(\mathcal{P}_{tot}A_{j})$. A value of $\mathcal{I}=1$ indicates a homogeneous dissipation, while values >1>1 and <1<1 indicate hot spots and cold spots, respectively. It can be observed, see Figure 12b, that the local power dissipation on the surface of the particles clusters in the same locations, where slip events are more common. For high degrees of eccentricity, more than ten times dissipation occurs on the major axis compared to the minor one. Accordingly, we propose the following microscopic mechanism: non-spherical particles establish a small number of strong contacts along their minor axis, where sliding is rare. To accommodate the linear velocity gradient, the particles exhibit a wobbling motion, resulting in the majority of power dissipation occurring at weaker contacts along the major axis. A more detailed microscopic analysis of the contribution of frictional forces to shear strength can be found in Appendix G.
@@ IV Summary and outlook
Our dissipation regime analysis revealed that flat, lentil-like particles exhibit an extended sliding regime compared to elongated, rice-like particles, even when the shape ratio is kept constant. This behavior was further clarified by examining the fluctuating energy components: oblate and prolate particles differ in how this energy is distributed and expressed in their dynamics. Moreover, slightly aspherical particles at low friction exhibit non-monotonic behavior, which we linked to enhanced rotational kinetic energy fluctuations.
@@
This work can be extended in several directions. First, the relationship between fabric anisotropy and stress transmission in non-spherical particle assemblies could be further examined within the framework proposed by Srivastava et al. [71]. Additionally, incorporating frictional effects into the anisotropic rheological model proposed by Nadler et al. [25] or the correlation lengths into extended kinetic theories [29] could enhance predictive capabilities for both natural and industrial applications.
Future work should explore the connection between the dissipation-based regimes identified in our study and the stress-based regimes (e.g., rate-dependent vs. rate-independent) discussed in the literature [13, 31, 53].
Another key avenue for investigation is the role of vortices and clusters in shear-driven particulate flows, as discussed in [69, 62, 41, 72]. The formation and dynamics of vortical structures could significantly impact shear banding and localized dissipation, necessitating a deeper physical understanding. Moreover, extending our analysis to dense suspensions [33] will help bridge the gap between granular and suspension rheology, elucidating the transition between friction-dominated and hydrodynamic-dominated flow regimes. Lubrication forces may become dominant at lower viscous numbers, especially for flat particles which enhance lubrication effects. Finally, a systematic study should be conducted on non-convex, potentially interlocking particles [73]. A deeper understanding of these interactions could inform the design of polycatenated materials [74] with novel particle architectures, ultimately improving control over their flow properties.
@@ Acknowledgements
We thank Professor Ali Ozel for sharing the preliminary input scripts on simple shear under quasi-static conditions, which contributed significantly to the work presented in [75]. We also acknowledge Lucas Fourel for expert assistance with visualization software and insightful advice on data presentation. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 945363. J.B. and J.-F.M. further appreciate the financial support from Bühler AG and the valuable discussions with Clément Zemerli and Richard Vonlanthen. M.T. acknowledges funding from the Swedish Research Council under grant No. 2021-04997.
@@
Data to support these findings is available on Zenodo [76]. Codes to run the simulations and the analysis are available on GitHub at https://github.com/JBil8/Simple_shear_granular_spheroids.git.
@@ Appendix A Contact interaction
$T_{H}=$ $2.94\left(5\sqrt{2}\pi\rho_{p}\over 4E^{*}\right)^{2/5}{a_{\text{min}}\over\dot{\gamma}d_{eq}}\,$ (23)
$T_{R}=$ ${\pi a_{\text{min}}\over 0.8766+0.163\nu}\sqrt{\rho_{p}\over G}\,$ (24)
where $G$ is the shear modulus and we assumed the typical impact velocity to be of the order of the one imposed by the shear flow ($\dot{\gamma}d_{eq}$), and $a_{\text{min}}$ is the smallest dimension of the smaller grains.
@@
$\text{subject to }G_{1}(\mathbf{x})=G_{2}(\mathbf{x})\,.$ (26)
@@ Appendix B Dissipation computation
$\mathcal{P}_{n}$ $=\sum_{N_{c}}\left[\frac{\mathbf{F}^{\gamma}_{n,i}}{2}\cdot\left(\mathbf{u}_{n,i}+\frac{1}{4m*}\mathbf{F}_{n,i}\right)+\right$.
$\quad\left.\frac{\mathbf{F}^{\gamma}_{n,i+1}}{2}\cdot\left(\mathbf{u}_{n,i+1/2}+\frac{1}{4m*}\mathbf{F}_{n,i+1}\right)\right]$
$\approx\sum_{N_{c}}\mathbf{F}_{t,i}\cdot\mathbf{u}_{t,i}\,$, (27)
@@
$\mathcal{P}_{n}=\sum_{N_{s}}\mathbf{F}_{t,i}\cdot\left(\frac{\bm{\delta}_{t,i}+\mathbf{F}_{t,i}/k_{t}}{\Delta t}\right)\approx\sum_{N_{s}}\mathbf{F}_{t,i}\cdot\mathbf{u}_{t,i}$ (28)
where the second equality is satisfied for small enough $\Delta t$ so that slip only occurs at the beginning of the time step and the tangential velocity can be considered constant during the interval. We verified both Equations 27 and 28 this by halving the time step size. In particular Equation 27 is tested for those simulations with the highest fluctuations $\mu_{p}=10$, $I=0.1$, corresponding to dominant normal dissipation, while Equation 28 is tested for $\mu_{p}=0.4$, $I=0.01$, where frictional dissipation is dominant. Figure 13 shows results within statistical error, implying that the chosen time step is sufficiently small.